OpenStudy (motherhen544):

Rewrite the expression without using a negative exponent. 3v^−2 Simplify your answer as much as possible.

OpenStudy (1davey29):

A negative exponent simply means to divide 1 by that term, so 3*v^-2 is the same as 3*(1/(v^2))=3/(v^2)

jhonyy9 (jhonyy9):

use formula a^-1 = 1/a

OpenStudy (motherhen544):

Can you put that in a picture because that looks really confusing

OpenStudy (motherhen544):

I just really don't get this concept at all

jhonyy9 (jhonyy9):

\[a^{-1}= 1/a\]

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (motherhen544):

Im still confused

OpenStudy (motherhen544):

like i said i just don't get it and I've never seen that formula before.

jhonyy9 (jhonyy9):

why ? there is y^(-2) = ?

OpenStudy (1davey29):

It's not really a formula, more of a rule of exponents. The concept uses rules of dividing exponents (ex. x^2/x=x^(2-1)=x), as 1=x^0 so 1/x=x^0/x^1=x^(0-1)=x^-1. Same with -2. x^0/x^2=x^(0-2)=x^-2

OpenStudy (motherhen544):

would it be 9v?

OpenStudy (motherhen544):

or no

OpenStudy (skullpatrol):

no

OpenStudy (motherhen544):

I don't get it at all

OpenStudy (1davey29):

The exponent only applies to the 3 if it's in parenthesis ((3v)^2). Otherwise, it just applies to the v.

OpenStudy (skullpatrol):

$$\huge 3v^{−2}=\dfrac3{v^2}$$

OpenStudy (motherhen544):

why did you put it in a fraction?

OpenStudy (skullpatrol):